# The fate of a single impurity in the Bose-Hubbard model

### v1

@@ Abstract
We map out the global phase diagram of a single mobile impurity in the two-dimensional Bose–Hubbard model, spanning the bath evolution from a compressible superfluid (SF) to an incompressible Mott insulator (MI) and the full range of impurity–bath coupling. Using sign-problem-free worm-algorithm quantum Monte Carlo method, we identify two sharply distinct localization mechanisms that organize the entire diagram. In the compressible SF, increasing impurity-bath coupling $|U_{\mathrm{ib}}|$ drives an interaction-driven winding-collapse crossover: a light, extended polaron with finite winding evolves continuously into a heavy polaron and ultimately a self-trapped state—a repulsive saturated bubble or an attractive bound cluster—even while the bath remains globally superfluid, establishing localization without any bath phase transition. By contrast, upon tuning the bath across the SF–MI transition at fixed impurity-bath coupling $U_{\mathrm{ib}}$, localization becomes compressibility controlled: the vanishing bath compressibility quenches polaronic dressing and collapses the impurity-centered density response, converting the polaron into an almost free defect and, deep in the MI, into a fully localized vacancy or particle defect. In the incompressible regime, localization proceeds via quantized defect formation, manifested by discrete changes in the total bath occupation. Together, our results provide a unified microscopic picture of impurity localization in correlated lattice bosons, governed by winding collapse in the SF and compressibility-driven localization across the SF-MI transition.
@@
Introduction– A single mobile impurity embedded in a quantum many-body medium provides a minimal yet stringent probe of correlations, coherence, and transport. In weakly interacting Bose gases, the impurity forms a well-defined Bose polaron: a coherent quasiparticle with finite residue, a renormalized effective mass, and an extended dressing cloud built from gapless density fluctuations[1, 2, 3, 3, 4, 5, 6, 7, 8, 9]. A central question is how this quasiparticle picture ultimately breaks down as correlations grow. In particular, can an impurity lose mobility and localize purely through interactions even in an otherwise uniform, translation-invariant superfluid, and how does this fate change when the bath itself becomes incompressible?
@@
The two-dimensional Bose–Hubbard model at unit filling ($\langle n_{\mathrm{b}}\rangle=1$) offers an ideal setting to address these questions [10, 11, 12, 13]. By tuning the intra-bath interaction $U_{\mathrm{b}}/t$, the bath evolves from a compressible superfluid (SF) to an incompressible Mott insulator (MI). Adding a single impurity with local impurity–bath coupling $U_{\mathrm{ib}}$ (repulsive or attractive) poses a sharply defined fate problem: what are the distinct regimes of impurity motion and localization across the full $(U_{\mathrm{b}}/t,U_{\mathrm{ib}}/t)$ parameter space, and what microscopic mechanisms separate them?
@@
While the limiting regimes are qualitatively understood [14, 15, 16, 17, 18, 19, 20], their global connection remains incomplete. In a compressible SF, most studies emphasize the coherent and delocalized polaron regime, where the impurity remains mobile and its dressing is supported by long-wavelength density fluctuations. In a deep MI, the bath has a gap and vanishing compressibility; weak coupling tends to leave the impurity as an almost free defect with minimal dressing, whereas strong coupling can bind the impurity to localized vacancy- or particle-defect. However, a unified picture of how impurity transport and dressing are lost—either by increasing $|U_{\mathrm{ib}}|$ in the SF or by quenching bath compressibility across the SF–MI transition—has not been established. In particular, interaction-driven winding-collapse localization without any bath phase transition in a uniform SF background has remained an open issue.
@@
Here we resolve this problem by constructing the global phase diagram of a single mobile impurity in the two-dimensional Bose–Hubbard model at $\langle n_{\mathrm{b}}\rangle=1$. We employ large-scale, sign-problem-free worm-algorithm quantum Monte Carlo [21, 22, 23, 24] to obtain unbiased equilibrium results across both SF and MI regimes. Crucially, we go beyond characterizing the impurity alone and track the microscopic bath response—the impurity-induced density redistribution and its collapse as the bath loses compressibility. Our diagnostics include impurity transport from winding statistics, impurity coherence from single-particle correlations, and the bath response from impurity-centered density modulations; in the MI, we further identify quantized vacancy/interstitial defects through discrete changes in the total bath occupation.
@@
The resulting phase diagram is organized by two sharply distinct localization mechanisms. Within the compressible SF, increasing $|U_{\mathrm{ib}}|$ drives an interaction-driven winding-collapse crossover: a light, extended polaron with finite winding evolves continuously into a heavy polaron and ultimately self-traps into a compact bound state—a repulsive saturated bubble or an attractive bound cluster—even while the bath remains globally superfluid. This establishes localization without any bath phase transition, controlled by the impurity–bath interaction through a loss of winding transport. Across the SF–MI transition at fixed $U_{\mathrm{ib}}$, localization becomes compressibility controlled: as bath compressibility is quenched, polaronic dressing and the impurity-induced density response collapse, and deep in the MI the impurity locks into a pinned vacancy (repulsive) or particle (attractive) defect with a quantized $\pm 1$ change in the total bath occupation. Together, these results provide a unified microscopic framework for impurity localization in correlated lattice bosons, governed by winding collapse in the SF and compressibility-driven localization across the SF-MI transition.
@@
Figure 1: Phase diagram of a single impurity in the two-dimensional Bose–Hubbard model in the $(U_{\mathrm{ib}}/t,\,U_{\mathrm{b}}/t)$ plane. The vertical dashed line separates repulsive ($U_{\mathrm{ib}}/t>0$) and attractive ($U_{\mathrm{ib}}/t<0$) impurity–bath couplings, while the horizontal dashed line marks the superfluid–Mott transition of the bath ($U_{\mathrm{b}}/t\simeq 16.7$). In the compressible superfluid region ($U_{\mathrm{b}}/t<16.7$), increasing $|U_{\mathrm{ib}}|/t$ drives an interaction-driven, winding-collapse polaron localization: a mobile light polaron continuously evolves into a heavy polaron and, for sufficiently strong coupling, into a saturated bubble (repulsive side) or a bound cluster (attractive side). In the incompressible Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 16.7$), the impurity localizes by quantized defect formation. Upon increasing $|U_{\mathrm{ib}}|/t$ beyond a threshold, the impurity changes abruptly from a nearly free defect to a pinned vacancy (repulsive) or particle (attractive) defect, signaled by an integer-valued bath occupation change. For intermediate couplings, the impurity remains only weakly dressed and essentially free. Along vertical cuts at fixed $U_{\mathrm{ib}}/t$, crossing the SF–MI boundary thus realizes compressibility-controlled localization. Black dots mark QMC data points; solid white lines are guides to the eye highlighting continuous crossover boundaries in the superfluid, while dotted lines indicate sharp transitions between distinct defect regimes in the Mott phase.
@@
Winding-number diagnostics– To characterize impurity mobility and the global coherence of the bath, we measure the winding numbers of both components. For the impurity, we record the winding number square of the impurity $\langle W_{\mathrm{imp}}^{2}\rangle=\langle W_{\mathrm{imp},x}^{2}+W_{\mathrm{imp},y}^{2}\rangle$, where $W_{\mathrm{imp},x}$ and $W_{\mathrm{imp},y}$ denote the net worldline crossings through the $x$ and $y$ directions. A finite $\langle W_{\mathrm{imp}}^{2}\rangle$ signals a mobile impurity with extended trajectories, while $\langle W_{\mathrm{imp}}^{2}\rangle\!\to\!0$ indicates complete localization. For the bath, the corresponding estimator gives the superfluid density $\rho_{\mathrm{s}}^{\mathrm{bath}}=\langle W_{{\mathrm{bath}},x}^{2}+W_{{\mathrm{bath}},y}^{2}\rangle/(2t_{b}\beta)$, directly related to the superfluid density via the Pollock–Ceperley relation [25].
@@
To remove lattice anisotropy, we perform a rotational averaging over all lattice sites with the same Euclidean distance $R=|\mathbf{r}|=\sqrt{r_{x}^{2}+r_{y}^{2}}$ (using the minimum–image convention), defining $C_{\mathrm{ib}}(R)=N_{R}^{-1}\!\sum_{|\mathbf{r}|=R}C_{\mathrm{ib}}(\mathbf{r})$, where $N_{R}$ counts sites satisfying $|\mathbf{r}|=R$. This smooth radial profile characterizes the impurity-centered bath density deformation: $C_{\mathrm{ib}}(R)<0$ signals a local depletion (bubble) for repulsive $U_{\mathrm{ib}}/t>0$, while $C_{\mathrm{ib}}(R)>0$ indicates a density accumulation (cluster) for attractive $U_{\mathrm{ib}}/t<0$. The cumulative excess $\Delta N(R)=\sum_{|\mathbf{r}|\leq R}C_{\mathrm{ib}}(\mathbf{r})$ gives the net bath particle-number change within a disk of radius $R$ centered at the impurity, whose amplitude and extent distinguish the light, heavy, and localized regimes discussed below.
@@
Together, these observations reveal a clear dichotomy in how impurity localizes in correlated Bose-Hubbard model, as shown in Fig. 1. In the compressible superfluid bath ($U_{\mathrm{b}}/t\lesssim 16.7$), localization is a continuous process: increasing $|U_{\mathrm{ib}}|/t$ gradually suppresses coherent winding and compresses the dressing cloud, producing the smooth crossover from light polarons to heavy polarons and ultimately to saturated-bubble (repulsive) or bound-cluster (attractive) states. We refer to this mechanism as interaction-driven winding-collapse polaron localization.
@@
In contrast, when impurity-bath coupling $U_{\mathrm{ib}}/t$ is held fixed and bath interaction $U_{\mathrm{b}}/t$ is increased across the SF–MI transition, the bath compressibility—not impurity-bath coupling strength—controls localization. Polaronic dressing collapses abruptly upon entering the incompressible regime, and the impurity instead undergoes an abrupt transition from a mobile nearly-free defect to a pinned localized defect carrying an integer excess (+1+1) or deficit (1-1) of bath occupation. We refer this as the compressibility-controlled localization mechanism.
@@
These two localization routes—continuous, interaction-driven winding-collapse in a superfluid bath and compressibility-controlled localization across the SF–MI transition, which at strong bath interaction (deep MI regiem) culminates in quantized defect formation—constitute two fundamentally distinct paradigms of impurity physics in strongly correlated bosonic matter.
@@
Figure 2: Impurity quasiparticle properties in a superfluid bath ($U_{\mathrm{b}}/t=13.3$). (a) Ground-state energy $E_{\mathrm{p}}/t$, (b) effective mass $m^{*}/m_{0}$ (with $m_{0}=1/2t=1/2$ the bare particle mass), and (c) quasiparticle residue $Z_{0}$ as functions of repulsive impurity-bath coupling $U_{\mathrm{ib}}/t$. All data correspond to finite impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle>0$, where a mobile light polaron is well defined. Increasing $U_{\mathrm{ib}}$ continuously enhances mass renormalization and suppresses $Z_{0}$, reflecting stronger impurity–bath correlations. Beyond $U_{\mathrm{ib}}/t\gtrsim 13.3$, the winding number vanishes (shown in Figure 3(e)), and the quasiparticle description breaks down, marking the boundary of the light-polaron regime. The ground-state energy changes sommthly throughout, confirming a continuous interaction-driven crossover rather than a sharp transition.
@@
Interaction-driven polaron localization in a superfluid bath: (i) Quasiparticle renormalization in the light-polaron regime. To characterize the impurity in a compressible superfluid background, we fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (deep in the SF phase) and increase the repulsive impurity-bath coupling $U_{\mathrm{ib}}/t>0$. Figure 2 shows the ground-state energy $E_{\mathrm{p}}/t$, effective mass $m^{*}/m_{0}$ with $m_{0}=1/2t=1/2$ the bare particle mass, and quasiparticle residue $Z_{0}$ extracted from imaginary-time Green’s functions, restricted to the regime with finite winding $\langle W_{\mathrm{imp}}^{2}\rangle>0$, where a mobile polaron is well defined.
@@
At weak impurity-bath coupling ($U_{\mathrm{ib}}/t\lesssim 10.0$), the impurity forms a light polaron: $m^{*}/m_{0}\lesssim 2.0$, $Z_{0}\gtrsim 0.5$, and the energy varies smoothly, indicating negligible dressing. For intermediate coupling ($10.0\lesssim U_{\mathrm{ib}}/t\lesssim 13.3$), $m^{*}/m_{0}$ increases while $Z_{0}$ decreases continuously, reflecting enhanced impurity–bath correlations and increasing mass renormalization, yet coherent transport persists. Beyond this range ($U_{\mathrm{ib}}/t\gtrsim 13.3$), the impurity winding number vanishes (shown in Fig. 3(e)), signaling the loss of coherent motion and the breakdown of the quasiparticle picture. Consequently, quantities such as $m^{*}/m_{0}$ increases abruplt and $Z_{0}$ decreses to zero, marking the boundary of the light-polaron to heavy-polaron regime. Throughout the mobile region, $E_{\mathrm{p}}/t$ evolves smoothly with $U_{\mathrm{ib}}/t$, confirming that the dressing process is a smooth crossover rather than a sharp transition.
@@
Figure 3: Real-space crossover from a light polaron to a saturated bubble in the superfluid bath ($U_{\mathrm{b}}/t=13.3$). (a) Cumulative bath-density response $\Delta N(R)$ for increasing impurity–bath coupling $U_{\mathrm{ib}}/t$. Red curves denote the light-polaron regime, the single orange curve highlights the heavy-polaron regime, and blue curves correspond to the saturated-bubble regime where the impurity becomes localized. (b) Impurity–centered correlation $C_{\mathrm{ib}}(R)$ and its on-site contrast $C_{\mathrm{ib}}(0)$ (inset). (c,d) Position $R_{\min}$ and depth $\Delta N_{\min}$ of the depletion minimum extracted from $\Delta N(R)$. (e) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$, exhibiting the continuous loss of coherent motion. (f) Bath superfluid density $\rho_{\mathrm{b}}$, which remains nearly constant across the crossover. The color evolution illustrates the continuous crossover from a mobile light polaron (red) to a compact heavy polaron (orange), and finally to a fully localized saturated bubble (blue).
@@
(ii) Real-space crossover from light polaron to heavy polaron and saturated bubble state. To elucidate how the impurity reshapes its surrounding superfluid environment, we fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (deep in the SF regime) and gradually increase the repulsive impurity–bath coupling $U_{\mathrm{ib}}/t>0$. Figure 3 summarizes the corresponding observables, revealing a continuous evolution from a mobile light polaron (red curves) to a heavy polaron (orange), and finally to a spatially saturated-bubble state (blue).
Figure 3(a) shows the cumulative bath-density response $\Delta N(R)$ for different impurity-bath coupling $U_{\mathrm{ib}}$. At weak coupling ($U_{\mathrm{ib}}/t<13.3$), $\Delta N(R)$ exhibits a shallow, broad minimum around $R_{\min}\!\approx\!4$–5 with $\Delta N_{\min}\lesssim-0.4$, indicating an extended depletion cloud, characteristic of a mobile light polaron. As $U_{\mathrm{ib}}/t$ increases to $13.3\!\lesssim\!U_{\mathrm{ib}}/t\!\lesssim\!20.0$, the depletion deepens and contracts to $R_{\min}\!\approx\!2$–3, while $\Delta N_{\min}$ approaches 0.8-0.8. However, the impurity winding number has already dropped close to zero, indicating that although the cloud remains spatially extended, coherent motion is lost. This intermediate regime corresponds to a heavy polaron: localized in real space but still surrounded by a finite-size many-body dressing cloud. For $U_{\mathrm{ib}}/t>20$, all $\Delta N(R)$ traces collapse onto a universal profile beyond $R\simeq 4$, with the depth approaching but not exactly reaching 1-1. This spatial convergence marks the onset of a saturated-bubble state: the depletion cloud has reached a fixed spatial extent and cannot shrink further, although its depth continues to increase with $U_{\mathrm{ib}}/t$. This saturation is spatial, not quantized—unlike the integer defects encountered in the incompressible Mott background, see the details in [26].
@@
Figure 3(b) shows the corresponding impurity-centered correlator $C_{\mathrm{ib}}(R)$, whose rapid on-site suppression (inset) tracks the same crossover. For $U_{\mathrm{ib}}/t>20$, $C_{\mathrm{ib}}(R>1)$ becomes nearly coupling-independent, confirming that the dressing cloud has reached its saturated spatial extent. Figure 3(c–f) quantify this continuum crossover. The extracted $R_{\min}$ and $\Delta N_{\min}$ clearly demonstrate the shrinking and deepening of the deformation cloud [Fig. 3(c,d)]. Meanwhile, the impurity winding estimator $\langle W_{\mathrm{imp}}^{2}\rangle$ [Fig. 3(e)] falls smoothly by nearly two orders of magnitude for $U_{\mathrm{ib}}/t>13.3$, signaling the loss of coherent worldline motion rather than a sharp phase transition from a light polaron to a heavy polaron. In contrast, the bath superfluid density $\rho_{\mathrm{b}}$ [(f)] remains finite and only weakly suppressed, confirming that the localization is interaction-driven while the bath superfluid stays globally coherent.

### v2

# Impurity Self-Trapping in Lattice Bose systems

@@ Abstract
We map out the global phase diagram of a single mobile impurity in the two-dimensional Bose–Hubbard model, spanning the bath evolution from a compressible superfluid (SF) to an incompressible Mott insulator (MI) and the full range of impurity–bath coupling. Using sign-problem-free worm-algorithm quantum Monte Carlo, we identify two distinct self-trapping mechanisms that organize the entire diagram. In the compressible SF, increasing impurity-bath coupling $|U_{\mathrm{ib}}|$ drives an interaction-driven self-trapping crossover signaled by a collapse of the impurity winding number: a light, extended polaron evolves continuously into a heavy polaron and ultimately into a self-trapped state—a repulsive saturated bubble or an attractive bound cluster—even while the bath remains globally superfluid, demonstrating self-trapping without any bath phase transition. By contrast, at fixed $U_{\mathrm{ib}}$ and upon tuning the bath across the SF–MI transition, self-trapping becomes compressibility controlled: the vanishing bath compressibility quenches polaronic dressing and collapses the impurity-centered density response, yielding a nearly free defect near criticality and, deep in the MI, a self-trapped vacancy- or particle-type defect bound to the impurity. In the incompressible regime, this proceeds via quantized defect formation, manifested by discrete changes in the total bath occupation. Together, our results provide a unified microscopic picture of impurity self-trapping in correlated lattice bosons, governed by winding collapse in the SF and by compressibility loss and defect quantization across the SF–MI boundary.
@@
Introduction– A single mobile impurity embedded in a quantum many-body medium provides a minimal yet stringent probe of correlations, coherence, and transport. In weakly interacting Bose gases, the impurity forms a well-defined Bose polaron: a coherent quasiparticle with finite residue, a renormalized effective mass, and an extended dressing cloud built from gapless density fluctuations[1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. A central question is how this quasiparticle picture breaks down as correlations grow. In particular, can an impurity lose mobility and self-trap purely through interactions even in an otherwise uniform, translation-invariant superfluid, and how does this fate change once the bath itself becomes incompressible?
@@
The two-dimensional Bose–Hubbard model at unit filling ($\langle n_{\mathrm{b}}\rangle=1$) provides an ideal setting to address these questions [11, 12, 13, 14]. By tuning the intra-bath interaction $U_{\mathrm{b}}/t$, the bath evolves from a compressible superfluid (SF) to an incompressible Mott insulator (MI). Adding a single impurity with a local impurity–bath coupling $U_{\mathrm{ib}}$ (repulsive or attractive) poses a sharply defined fate problem: what are the distinct regimes of impurity motion and self-trapping across the full $(U_{\mathrm{b}}/t,U_{\mathrm{ib}}/t)$ parameter space, and what microscopic mechanisms separate them?
@@
While the limiting regimes are qualitatively understood [15, 16, 17, 18, 19, 20, 21], their global connection has remained incomplete. In the compressible SF, most studies emphasize the coherent polaron regime, where the impurity remains mobile and its dressing is supported by long-wavelength density fluctuations. In the deep MI, the bath has a gap and vanishing compressibility; at weak coupling the impurity is nearly a free defect with minimal dressing, whereas at strong coupling it can bind to and pin a vacancy- or particle-type defect. However, a unified microscopic account of how impurity transport and dressing are lost—either by increasing $|U_{\mathrm{ib}}|$ within the SF or by quenching bath compressibility across the SF–MI transition—is still lacking. In particular, interaction-driven self-trapping in a uniform SF background without any bath phase transition has not been established in a controlled, unbiased setting.
@@
Here we resolve this problem by constructing the global phase diagram of a single mobile impurity in the two-dimensional Bose–Hubbard model at $\langle n_{\mathrm{b}}\rangle=1$. We employ large-scale, sign-problem-free worm-algorithm quantum Monte Carlo [22, 23, 24, 25] to obtain unbiased equilibrium results across both SF and MI regimes. Crucially, we go beyond characterizing the impurity alone and track the microscopic bath response—the impurity-induced density redistribution and its collapse as the bath loses compressibility. Our diagnostics quantify impurity transport via impurity winding statistics, impurity coherence via single-particle correlations, and the bath response via impurity-centered density modulations; in the MI we further identify quantized vacancy/particle defects through discrete changes in the total bath occupation.
@@
The resulting phase diagram is organized by two sharply distinct self-trapping mechanisms. Within the compressible SF, increasing $|U_{\mathrm{ib}}|$ drives an interaction-driven winding-collapse crossover: a light, extended polaron with finite impurity winding evolves continuously into a heavy polaron and ultimately self-traps into a compact bound state—a repulsive saturated bubble or an attractive bound cluster—even while the bath remains globally superfluid. This demonstrates self-trapping without any bath phase transition, controlled by the impurity–bath interaction through a collapse of winding transport. Across the SF–MI transition at fixed $U_{\mathrm{ib}}$, self-trapping becomes compressibility controlled: as the bath compressibility is quenched, polaronic dressing and the impurity-induced density response collapse, and deep in the MI the impurity locks to a pinned vacancy (repulsive) or particle (attractive) defect with a quantized $\pm 1$ change in the total bath occupation. Together, these results provide a unified microscopic framework for impurity self-trapping in correlated lattice bosons, governed by winding collapse in the SF and by compressibility loss and defect quantization across the SF–MI boundary.
@@
Figure 1: Phase diagram of a single impurity in the two-dimensional Bose–Hubbard model in the $(U_{\mathrm{ib}}/t,\,U_{\mathrm{b}}/t)$ plane. The vertical dashed line separates repulsive ($U_{\mathrm{ib}}/t>0$) and attractive ($U_{\mathrm{ib}}/t<0$) impurity–bath couplings, while the horizontal dashed line marks the superfluid–Mott transition of the bath ($U_{\mathrm{b}}/t\simeq 16.7$). In the compressible superfluid region ($U_{\mathrm{b}}/t<16.7$), increasing $|U_{\mathrm{ib}}|/t$ drives an interaction-driven self-trapping crossover signaled by a collapse of the impurity winding number: a mobile light polaron evolves continuously into a heavy polaron and, at strong coupling, into a saturated bubble (repulsive side) or a bound cluster (attractive side). In the incompressible Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 16.7$), self-trapping proceeds via quantized defect formation. Beyond a threshold in $|U_{\mathrm{ib}}|/t$, the impurity changes abruptly from a nearly free defect to a pinned vacancy (repulsive) or particle (attractive) defect, signaled by an integer jump $\Delta N_{\mathrm{b}}=\mp 1$ in the total bath occupation. For intermediate couplings, the impurity remains weakly dressed and essentially free. Along vertical cuts at fixed $U_{\mathrm{ib}}/t$, crossing the SF–MI boundary thus realizes compressibility-controlled self-trapping. Black dots mark QMC data points; solid white lines are guides to the eye highlighting continuous crossover boundaries in the superfluid, while dotted lines indicate sharp transitions between distinct defect regimes in the Mott phase.
@@
Winding-number diagnostics– To characterize impurity mobility and the global coherence of the bath, we measure the winding numbers of both components. For the impurity, we record the winding number square of the impurity $\langle W_{\mathrm{imp}}^{2}\rangle=\langle W_{\mathrm{imp},x}^{2}+W_{\mathrm{imp},y}^{2}\rangle$, where $W_{\mathrm{imp},x}$ and $W_{\mathrm{imp},y}$ denote the net worldline crossings through the $x$ and $y$ directions. A finite $\langle W_{\mathrm{imp}}^{2}\rangle$ signals a mobile impurity with extended trajectories, whereas a strong suppression of $\langle W_{\mathrm{imp}}^{2}\rangle$ toward zero signals an effectively self-trapped impurity. For the bath, the corresponding estimator gives the superfluid density $\rho_{\mathrm{b}}=\langle W_{{\mathrm{b}},x}^{2}+W_{{\mathrm{b}},y}^{2}\rangle/(2t_{\rm b}\beta)$, directly related to the superfluid density via the Pollock–Ceperley relation [26].
@@
To remove lattice anisotropy, we perform a rotational averaging over all lattice sites with the same Euclidean distance $R=|\mathbf{r}|=\sqrt{r_{x}^{2}+r_{y}^{2}}$ (using the minimum–image convention), defining $C_{\mathrm{ib}}(R)=N_{R}^{-1}\!\sum_{|\mathbf{r}|=R}C_{\mathrm{ib}}(\mathbf{r})$, where $N_{R}$ counts sites satisfying $|\mathbf{r}|=R$. This smooth radial profile characterizes the impurity-centered bath density deformation: $C_{\mathrm{ib}}(R)<0$ signals a local depletion (bubble) for repulsive $U_{\mathrm{ib}}/t>0$, while $C_{\mathrm{ib}}(R)>0$ indicates a density accumulation (cluster) for attractive $U_{\mathrm{ib}}/t<0$. The cumulative excess $\Delta N(R)=\sum_{|\mathbf{r}|\leq R}C_{\mathrm{ib}}(\mathbf{r})$ gives the net bath particle-number change within a disk of radius $R$ centered at the impurity, whose amplitude and extent distinguish the light, heavy, and self-trapped regimes discussed below.
@@
Together, these observations reveal a clear dichotomy in how a single impurity self-traps in the two-dimensional Bose–Hubbard model, as summarized in Fig. 1. In the compressible superfluid bath ($U_{\mathrm{b}}/t\lesssim 16.7$), self-trapping is continuous: increasing $|U_{\mathrm{ib}}|/t$ progressively suppresses coherent impurity winding and compresses the dressing cloud, yielding a smooth crossover from light to heavy polarons and, at strong coupling, to a saturated bubble (repulsive) or a bound cluster (attractive) state. We refer to this route as interaction-driven, winding-collapse self-trapping.
@@
In contrast, along vertical cuts at fixed $U_{\mathrm{ib}}/t$, increasing $U_{\mathrm{b}}/t$ across the SF–MI transition renders self-trapping compressibility controlled: as the bath compressibility is quenched, polaronic dressing collapses upon entering the incompressible regime, and the impurity crosses over from a mobile, nearly free defect to a pinned defect bound to an integer bath-occupation change, $\Delta N_{\mathrm{b}}=\pm 1$ (vacancy/particle for repulsive/attractive $U_{\mathrm{ib}}$). We refer to this as compressibility-controlled self-trapping.
@@
These two routes—a continuous, interaction-driven winding-collapse mechanism in the superfluid and a compressibility-controlled mechanism tied to quantized defect formation in the Mott insulator—define two fundamentally distinct paradigms of impurity self-trapping in strongly correlated lattice bosons.
@@
Figure 2: Impurity quasiparticle properties in a superfluid bath ($U_{\mathrm{b}}/t=13.3$). (a) Ground-state energy $E_{\mathrm{p}}/t$, (b) effective-mass ratio $m^{*}/m_{0}$, and (c) quasiparticle residue $Z_{0}$ versus the repulsive impurity–bath coupling $U_{\mathrm{ib}}/t$. Here $m_{0}$ is the bare lattice mass of the impurity (nearest-neighbor hopping). The quasiparticle diagnostics $m^{*}$ and $Z_{0}$ are shown in the mobile regime with finite impurity winding, $\langle W_{\mathrm{imp}}^{2}\rangle>0$. Increasing $U_{\mathrm{ib}}$ continuously enhances mass renormalization and suppresses $Z_{0}$, signaling stronger impurity–bath correlations. At $U_{\mathrm{ib}}/t\gtrsim 13.3$, the impurity winding collapses to zero [Fig. 3(e)], and the quasiparticle description breaks down, marking the boundary of the light-polaron regime. The smooth evolution of $E_{\mathrm{p}}/t$ indicates a continuous interaction-driven crossover rather than a sharp transition.
@@
Interaction-driven polaron self-trapping in a superfluid bath: (i) Quasiparticle renormalization in the light-polaron regime. To characterize the impurity in a compressible superfluid background, we fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (deep in the SF phase) and increase the repulsive impurity–bath coupling $U_{\mathrm{ib}}/t>0$. Figure 2 shows the ground-state energy $E_{\mathrm{p}}/t$, the effective-mass ratio $m^{*}/m_{0}$, and the quasiparticle residue $Z_{0}$ extracted from imaginary-time Green’s functions. The quasiparticle quantities are reported in the mobile regime with finite impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle>0$, where a coherent polaron is well defined.
@@
At weak coupling ($U_{\mathrm{ib}}/t\lesssim 10.0$), the impurity forms a light polaron: $m^{*}/m_{0}\lesssim 2.0$ and $Z_{0}\gtrsim 0.5$, with $E_{\mathrm{p}}/t$ varying smoothly, consistent with relatively weak dressing. For intermediate coupling ($10.0\lesssim U_{\mathrm{ib}}/t\lesssim 13.3$), $m^{*}/m_{0}$ increases and $Z_{0}$ decreases continuously, reflecting enhanced impurity–bath correlations while coherent transport persists. Beyond $U_{\mathrm{ib}}/t\gtrsim 13.3$, the impurity winding collapses to zero [Fig. 3(e)], signaling the loss of coherent motion and the breakdown of the quasiparticle picture: $m^{*}$ rapidly grows (effectively diverges as $\langle W_{\mathrm{imp}}^{2}\rangle\!\to\!0$) and $Z_{0}$ drops toward zero. Throughout the mobile regime, $E_{\mathrm{p}}/t$ evolves smoothly with $U_{\mathrm{ib}}/t$, confirming that the dressing and eventual self-trapping develop via a continuous crossover rather than a sharp transition.
@@
Figure 3: Real-space crossover from a light polaron to a saturated bubble in a superfluid bath ($U_{\mathrm{b}}/t=13.3$). (a) Cumulative bath-density response $\Delta N(R)$ for increasing repulsive impurity–bath coupling $U_{\mathrm{ib}}/t$. Red curves denote the light-polaron regime, the orange curve highlights an intermediate heavy-polaron regime, and blue curves correspond to the saturated-bubble regime where the impurity is self-trapped. (b) Impurity-centered correlator $C_{\mathrm{ib}}(R)$ and its on-site value $C_{\mathrm{ib}}(0)$ (inset). (c,d) Position $R_{\min}$ and depth $\Delta N_{\min}$ of the depletion minimum extracted from $\Delta N(R)$. (e) Impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$, showing a continuous loss of coherent motion. (f) Bath superfluid density $\rho_{\mathrm{b}}$, which remains nearly unchanged across the crossover. The color progression illustrates a continuous evolution from a mobile light polaron (red) to an immobile but still dressed heavy polaron (orange), and finally to a fully self-trapped saturated bubble (blue).
@@
(ii) Real-space crossover from light polaron to heavy polaron and saturated bubble. To elucidate how the impurity reshapes its surrounding superfluid environment, we fix the bath interaction at $U_{\mathrm{b}}/t=13.3$ (deep in the SF regime) and increase the repulsive impurity–bath coupling $U_{\mathrm{ib}}/t>0$. Figure 3 summarizes the corresponding observables and reveals a continuous evolution from a mobile light polaron (red curves), to an heavy polaron (orange), and finally to a spatially saturated bubble (blue).
Figure 3(a) shows the cumulative bath-density response $\Delta N(R)$. At weak coupling ($U_{\mathrm{ib}}/t\lesssim 13.3$), $\Delta N(R)$ exhibits a shallow, broad minimum around $R_{\min}\approx 4$–5 with $\Delta N_{\min}\gtrsim-0.4$, consistent with an extended depletion cloud accompanying a mobile light polaron. Upon increasing $U_{\mathrm{ib}}/t$ into the intermediate range ($13.3\lesssim U_{\mathrm{ib}}/t\lesssim 20$), the deformation deepens and contracts to $R_{\min}\approx 2$–3 while $\Delta N_{\min}$ approaches 0.8-0.8. At the same time, the impurity winding $\langle W_{\mathrm{imp}}^{2}\rangle$ drops to (nearly) zero [Fig. 3(e)], indicating that coherent transport is lost even though a finite-range dressing cloud persists. We identify this as a heavy-polaron regime: the impurity is effectively immobile, yet still surrounded by a compact many-body deformation.
For stronger coupling ($U_{\mathrm{ib}}/t\gtrsim 20$), the $\Delta N(R)$ curves become nearly coupling independent for $R\gtrsim 3$–4, approaching a common spatial profile whose extent no longer shrinks appreciably with $U_{\mathrm{ib}}/t$. This marks the onset of a saturated-bubble regime: the depletion cloud has reached a saturated spatial size, while its depth continues to evolve with $U_{\mathrm{ib}}/t$. Importantly, this saturation is not quantized (unlike the integer defects in the incompressible Mott background; see [27]).
@@
Figure 3(b) shows the impurity-centered correlator $C_{\mathrm{ib}}(R)$, whose suppression at short distances (inset) tracks the same crossover. For $U_{\mathrm{ib}}/t\gtrsim 20$, $C_{\mathrm{ib}}(R\!>\!1)$ becomes weakly dependent on $U_{\mathrm{ib}}/t$, corroborating that the dressing cloud has reached a saturated spatial extent. Panels (c,d) quantify the contraction and deepening of the deformation cloud through $R_{\min}$ and $\Delta N_{\min}$. Meanwhile, $\langle W_{\mathrm{imp}}^{2}\rangle$ decreases smoothly by nearly two orders of magnitude across the intermediate regime [Fig. 3(e)], consistent with a continuous crossover rather than a sharp phase transition. In contrast, the bath superfluid density $\rho_{\mathrm{b}}$ remains finite and is only weakly affected [Fig. 3(f)], confirming that the impurity self-traps through interaction-driven winding collapse while the bath stays globally superfluid.